# LF-Steering: Latent Feature Activation Steering for Enhancing Semantic Consistency in Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) often generate inconsistent responses when prompted with semantically equivalent paraphrased inputs. Recently, activation steering, a technique that modulates LLM behavior by adjusting their latent representations during inference time, has been explored to improve the semantic consistency of LLMs. However, these methods typically operate at the model component level, such as layer hidden states or attention heads. They face a challenge due to the “polysemanticity issue”, where the model components of LLMs typically encode multiple entangled features, making precise steering difficult. To address this challenge, we drill down to feature-level representations and propose LF-Steering, a novel activation steering approach to precisely identify latent feature representations responsible for semantic inconsistency. More specifically, our method maps the hidden states of relevant transformer layer into a sparsely activated, high-dimensional feature space based on a sparse autoencoder (SAE), ensuring model steering based on decoupled feature representations with minimal interference. Comprehensive experiments on both NLU and NLG datasets demonstrate the effectiveness of our method in enhancing semantic consistency, resulting in significant performance gains for various NLU and NLG tasks.
Figure 1: Semantic inconsistency problem in LLMs and the grey-box approache for addressing this problem.
@@ 1 Introduction
Existing semantic consistency enhancing approaches for LLMs can be broadly classified into two categories: end-to-end “black-box” methods and interpretability-oriented “grey-box” approaches with greater transparency (Figure 1). In “black-box” methods, data-driven supervised fine-tuning (SFT) (Ouyang et al., 2022; Fierro et al., 2024; Zhao et al., 2024) is commonly employed, using pairs of prompts and model responses with semantically equivalent meanings as training data. While these methods have shown effectiveness, they come with significant costs in data engineering and LLM fine-tuning. Recently, Yang et al. (2024) propose a “grey-box” method aimed at improving the semantic consistency of LLMs from a more transparent perspective. This approach identifies specific internal representations (e.g., attention head outputs) responsible for semantic inconsistency, and subsequently adjusts them toward a more semantically consistent direction to steer model behavior. However, due to the “polysemanticity issue” (Elhage et al., 2022), individual model components (i.e., neurons, attention heads, etc.) of LLMs typically encode a mixture of multiple unrelated features, making it challenging to precisely adjust model behaviors based on these entangled representations.
To address this challenge, we drill down to finer-grained feature-level representations of LLMs and propose a novel approach named “LF-Steering” to precisely identify and modify the relevant latent feature representations responsible for semantic inconsistency. The features are constructed by decomposing transformer layer representations into a sparsely activated, higher-dimensional feature space. This ensures that the model steering is based on decoupled feature representations with minimal interference. Our approach involves the following key steps. First, we apply a sparse autoencoder (SAE) to map the hidden states of the transformer layers into a feature space, and then identify relevant features that significantly impact on the LLM’s semantic consistency. After that, we selectively modify the top-k identified features to improve the model’s semantic consistency.
@@
  • Our proposed method is a generic activation steering technique applicable across a wide range of tasks. Extensive experiments demonstrate its effectiveness, achieving SOTA performance over existing leading activation steering methods in enhancing semantic consistency of LLMs. It outperforms the baselines in prediction accuracy by up to 8.9% on NLU datasets. For NLG datasets, accuracy can be improved by up to 7.12%.
@@ 2 Related Work
Semantic Consistency. LLMs tend to produce semantic inconsistent results (Gan and Mori, 2023; Fierro et al., 2024; Wang et al., 2024) when prompted with semantically equivalent paraphrased inputs. Current methods for improving semantic consistency of LLMs fall into two categories: end-to-end “black-box” methods and interpretability-oriented “grey-box” approaches. “Black-box” methods primarily relied on expert-designed prompting strategies (Raj et al., 2023) or SFT techniques (Ouyang et al., 2022; Zhou et al., 2022; Fierro et al., 2024; Zhao et al., 2024). Among them, Raj et al. (2023) leveraged an Ask-to-Choose (A2C) prompting technique to improve the accuracy as well as the semantic consistency of LLMs. Additionally, Zhao et al. (2024) proposed a data augmentation method to generate synthetic data consisting of pairs of paraphrased prompts and LLM responses for SFT, aiming to achieve the same end. Recently, Yang et al. (2024) attempted to address this issue through a more transparent “grey-box” method. This method involves identifying specific internal hidden state representations (such as attention head outputs) contributing to semantic inconsistency and subsequently adjusting these representations to achieve greater semantic consistency. Despite its effectiveness, the method was limited to operating on model components subjective to the aforementioned “polysemanticity issue”.
@@
Activation Steering. Activation steering techniques are emerged to modify specific model behaviors by adjusting their internal representations (i.e., the activation values of neurons). Existing activation steering approaches mostly rely on the hidden states of transformer layers. For instance, Turner et al. (2023) introduced an activation addition approach, which computes a steering vector from contrastive samples to guide model behaviors, such as generating specific sentiments or topics in its responses. Rimsky et al. (2024) employed mass mean activation difference to enhance model steering performance. Other methods operated on the outputs of attention heads to steer model behaviors. For example, Li et al. (2023) utilized linear probing to identify the attention heads correlated with truthful responses. By adjusting the outputs of these attention heads, they enhanced the truthfulness of an LLM’s responses. Chen et al. (2024) proposed a method to enhance the truthfulness of LLMs’s responses by modifying hidden state outputs along multiple orthogonal truthful directions identified by a linear probe.
@@
Sparse Autoencoder. The study of sparse autoencoders (SAEs) was first introduced by Agarwal et al. (2016) as an unsupervised learning method for feature extraction. More recently, SAEs have been used to disentangle representations within LLMs for analyzing polysemanticity phenomena in these models (Bricken et al., 2023). For example, Cunningham et al. (2023) trained SAEs with L1 loss to map the internal activations of LLMs into a higher-dimensional feature space, using these feature representations to interpret model behaviors. More recently, gated sparse autoencoder (Rajamanoharan et al., 2024) was proposed to address the in-balance between accurate reconstruction and activation sparsity caused by L1 loss biases. Gao et al. (2024) further proposed TopK SAE, leveraging a Top-K activation function to enforce sparsity constraints more effectively to this end.
In this paper, we adopt a TopK SAE to decompose transformer layer representations into a sparsely activated, higher-dimensional feature space, ensuring that the features are decoupled with minimal interference for precise steering.
Figure 2: The flowchart of our method. (1) Consistency Relevant Features Locating consists two steps. We first use the method proposed by (Yang et al., 2024) to locate the top-1 transformer layer and then pretrain a SAE and use it to locate the features responsible for the semantic inconsistencies in the LLM. (2) Steering the LLM towards greater semantic consistency by adjusting the identified top-K features.
@@ 3 Methodology
As illustrated in Figure 2, our proposed method consists of two main steps: Consistency Relevant Features Locating and Consistency Feature Steering. In the first step, we employ a coarse-to-fine locating strategy. Initially, we use the method proposed by (Yang et al., 2024) to identify the top-1 transformer layer that has the most significant impact on the model’s semantic inconsistency. Subsequently, we pretrain a spare autoencoder (SAE) and use it to project the hidden states of the identified transformer layer into a higher-dimensional feature space. Within this feature space, we locate the top-k features that are most relevant to the LLM’s semantic inconsistency. In the second step, we steer the LLM towards generating more semantically consistent responses by adjusting the identified top-k features.
@@ 3.1 Consistency Relevant Features Locating
Top-1 Transformer Layer Locating. We use the method proposed by (Yang et al., 2024) to locate the top-1 transformer layer influences the LLM’s semantic consistency. The idea behind this method is that the layer behaving similarly to the target LLM contribute to semantic inconsistency issues. Conversely, layers showing significant behavioral differences are less likely to be the cause. Based on this assumption, the layer locating dataset $\mathcal{D}_{l}$ was first built, comprising consistent pairs in the form $([m,n],l)$, where $m$ is the input prompt and $n$ is a rephrased version of $m$, which can be generated using large-scale LLMs like GPT-4111https://platform.openai.com/docs/api-reference/chat. The variable $l$ denotes whether the LLM outputs are consistent for $m$ and $n$, which can also be evaluated by large-scale LLMs. Following (Yang et al., 2024), the $\mathcal{D}_{l}$ dataset, consisting of 500 instances, was randomly split into a training set and a test set with a 4:1 ratio. Next, for each transformer layer, we train a classifier that takes the hidden states of the corresponding layer for prompts $m$ and $n$ as inputs and uses the consistency label $l$ as the ground truth. High classifier accuracy on the test set indicates that the layer closely mirrors the behavior of the LLM, while low accuracy suggests that the layer is less relevant to semantic inconsistency. The top-1 performing transformer layer is identified by ranking the classification accuracy.
TopK SAE Pretraining. Since the hidden states of the transformer layers contain entangled feature representations (Elhage et al., 2022), this hinders the precise locating of features related to semantic consistency. To address this, it is essential to delve into finer-grained feature-level representations for accurate features identification. To achieve this, we pretrain a sparse autoencoder (SAE) to decompose the transformer layer representations into a higher-dimensional, sparsely activated feature space, resulting in decoupled feature-level representations. In specific, we pretrain a TopK SAE (Gao et al., 2024) by using the representations from the identified top-1 transformer layer, leveraging 850K public data $\mathcal{D}_{o}$, collected from various domains  [^2] https://huggingface.co/datasets/togethercomputer/RedPajama-Data-1T-Sample. Rather than using task-specific data, this choice ensures that the learned feature representations are generalizable. More specifically, the TopK SAE consists of two main components: the SAE encoder and the SAE decoder. The encoder maps the model’s hidden states into a high-dimensional feature space and applies the TopK activation function to retain the $k$ most significant latent features, ensuring sparse activation. The decoder then reconstructs the original hidden states from these top $k$ latent features. The training process is formally represented as follows:
$\begin{gathered}z(d)=\text{TopK}(W_{\text{enc}}(h_{\text{top-1}}(d)-b_{\text{pre}}))\\ \hat{h}_{\text{top-1}}(d)=W_{\text{dec}}z(d)+b_{\text{pre}},\\ \mathcal{L}=\frac{1}{|\mathcal{D}_{d}|}\sum_{d}\|h_{\text{top-1}}(d)-\hat{h}_{\text{top-1}}(d)\|_{2}^{2},d\in\mathcal{D}_{o}.\\ \end{gathered}$ (1)
@@
Here, $d$ is the data instance from the $\mathcal{D}_{o}$. $h_{\text{top-1}}(d)$ and $\hat{h}_{\text{top-1}}(d)$ denote the hidden states and the reconstructed hidden states from the identified top-1 transformer layer respectively. $W_{\text{enc}}$, $W_{\text{dec}}$, and $b_{\text{pre}}$ represent the encoder weights, decoder weights, and the corresponding bias term, respectively. The $\mathcal{L}$ is the training loss.
@@
The TopK SAE (Gao et al., 2024) used in our paper decomposes the 4,096 dimensional hidden states to 65,536 dimensional features. We select the top 192 latent features to recover the original hidden states. The training process is performed with a batch size of 1 per GPU, utilizing 8xV100 GPUs each with 32 GB of memory. We train the TopK SAE for a total of 32,000 steps with the learning rate is set to 2e-4.
@@
Top-K Features Locating. To identify features related to semantic consistency within the feature space decomposed from the pretrained SAE, we build dataset $\mathcal{D}_{\text{f}}$ consists of 500 contrastive examples in the form $([u,v])$, where $u$ represents the input prompt and $v$ is its rephrased version obtained from GPT-4. The key difference is that $u$ generates a correct result when processed by the LLM, whereas $v$ produces an erroneous response from the target LLM. The correctness labels are assigned based on whether their predicted outputs are identical for NLU tasks. For NLG tasks, GPT-4 is also used to evaluate the correctness label.
@@
Next, we input the contrastive examples $u$ and $v$ from $D_{\text{f}}$ into the LLM to obtain their respective last token feature representations from the pretrained SAE. We then calculate the differences between these feature representations and select the top k positions with the largest feature differences as the identified key features related to LLM’s semantic inconsistency. This process is formally represented as follows:
@@
Here, $z(u)$ and $z(v)$ are the last token feature activations of $u$ and $v$ respectively. $g$ is the average feature differences across all feature positions, and $|g|$ is the size of $g$. $\mathcal{I}$ contains the indexes of the identified top k features. $t$ is a predefined feature difference threshold.
@@ 3.2 Consistency Feature Steering
In this step, we make precise adjustments to the identified feature positions to enhance semantic consistency. Specifically, when an input prompt is provided during inference, we check whether its activated feature positions from the pretrained SAE are present in the identified top-K feature indexes from last step. If so, we add the corresponding consistency feature bias (average feature differences) on it to guide the LLM towards greater semantic consistency. Formally, this steering process is calculated as follows:
@@
Here, $q$ is the given input prompt, $z(q)$ is the activated features from the pretrained SAE. $\mathds{1}$ denotes the indicator function of feature positions $i$ for steering. $b_{i}$ is the consistency feature bias. $\alpha$ is a hyperparameter that controls the strength of the steering. In the subsequent experiments, we set $\alpha$ to a value of $20$.

### v2

@@ Abstract
Large Language Models (LLMs) often generate inconsistent responses when prompted with semantically equivalent paraphrased inputs. Recently, activation steering, a technique that modulates LLMs’ behaviours by adjusting their latent representations during inference time, has been explored to improve the semantic consistency of LLMs. However, these methods typically operate at the model component level, such as layer hidden states or attention head outputs. They face a challenge due to the “polysemanticity issue”, where the model components of LLMs typically encode multiple entangled features, making precise steering difficult. To address this challenge, we drill down to feature-level representations and propose LF-Steering, a novel activation steering approach to precisely identify latent feature representations responsible for semantic inconsistency. More specifically, our method maps the hidden states of the relevant transformer layer into a sparsely activated, high-dimensional feature space based on a sparse autoencoder (SAE), ensuring model steering based on decoupled feature representations with minimal interference. Comprehensive experiments on NLU and NLG datasets demonstrate the effectiveness of our method in enhancing semantic consistency, resulting in significant performance gains for various NLU and NLG tasks.
Figure 1: Semantic inconsistency in LLMs and how a “grey-box” approach based on activation steering addresses this problem.
@@ 1 Introduction
Existing approaches to improve semantic consistency for LLMs can be broadly classified into two categories: end-to-end “black-box” methods and interpretability-oriented “grey-box” approaches with greater transparency (Figure 1). In “black-box” methods, data-driven supervised fine-tuning (SFT) (Ouyang et al., 2022; Fierro et al., 2024; Zhao et al., 2024) is commonly employed, using pairs of prompts and model responses with semantically equivalent meanings as training data. While these methods have shown effectiveness, they come with significant costs in data engineering and LLM fine-tuning. Recently, Yang et al. (2024) propose a “grey-box” method aimed at improving the semantic consistency of LLMs from a more transparent perspective. This approach identifies specific internal representations (e.g., attention head outputs) responsible for semantic inconsistency, and subsequently adjusts them toward a more semantically consistent direction to steer model behaviour. However, due to the “polysemanticity issue” (Elhage et al., 2022), individual model components (i.e., neurons, attention heads, etc.) of an LLM typically encode a mixture of multiple unrelated features, making it challenging to precisely adjust model behaviours based on these entangled representations.
To address this challenge, we drill down to finer-grained feature-level representations of LLMs and propose a novel approach named “LF-Steering” to precisely identify and modify the relevant latent feature representations responsible for semantic inconsistency. The features are constructed by decomposing transformer layer representations into a sparsely activated, higher-dimensional feature space. This ensures that the model steering is based on decoupled feature representations with minimal interference. Our approach involves the following key steps. First, we apply a sparse autoencoder (SAE) to map the hidden states of the transformer layers into a feature space, and then identify relevant features that significantly impact on the LLM’s semantic consistency. After that, we selectively modify the identified key features to improve the model’s semantic consistency.
@@
  • Our proposed method is a generic activation steering technique to address semantic inconsistency across a wide range of tasks. Extensive experiments demonstrate its effectiveness, achieving SOTA performance over existing leading activation steering methods in enhancing semantic consistency of LLMs. It outperforms the baselines in prediction accuracy by up to 8.9% on NLU datasets. For NLG datasets, accuracy can be improved by up to 7.12%.
@@ 2 Related Work
Semantic Consistency. LLMs tend to produce semantic inconsistent results (Gan and Mori, 2023; Fierro et al., 2024; Wang et al., 2024) when prompted with semantically equivalent paraphrased inputs. Current methods for improving semantic consistency of LLMs fall into two categories: end-to-end “black-box” methods and interpretability-oriented “grey-box” approaches. “Black-box” methods primarily relied on expert-designed prompting strategies (Raj et al., 2023) or SFT techniques (Ouyang et al., 2022; Zhou et al., 2022; Fierro et al., 2024; Zhao et al., 2024). Among them, Raj et al. (2023) leveraged an Ask-to-Choose (A2C) prompting technique to improve the accuracy as well as the semantic consistency of LLMs. Additionally, Zhao et al. (2024) proposed a data augmentation method to generate synthetic data consisting of pairs of paraphrased prompts and LLM responses for SFT, aiming to achieve the same end. Recently, Yang et al. (2024) attempted to address this issue through a more transparent “grey-box” method. Yang et al. (2024) applied an activation steering technique to identify specific internal hidden state representations (such as attention head outputs) contributing to semantic inconsistency and subsequently adjusted these representations to achieve greater semantic consistency. Despite its effectiveness, the method was limited to operating on model components subjective to the aforementioned “polysemanticity issue”.
@@
Activation Steering. Activation steering techniques are emerged to modify specific model behaviors by adjusting their internal representations (i.e., the activation values of neurons). Existing activation steering approaches mostly rely on the hidden states of transformer layers. For instance, Turner et al. (2023) introduced an activation addition approach, which computes a steering vector from contrastive samples to guide model behaviours, such as generating specific sentiments or topics in its responses. Rimsky et al. (2024) employed mass mean activation difference to improve model steering performance. Other methods operated on the outputs of attention heads to steer model behaviors. For example, Li et al. (2023) utilised linear probing to identify the attention heads correlated with truthful responses. By adjusting the outputs of these attention heads, they enhanced the truthfulness of an LLM’s responses. Chen et al. (2024) proposed a method to enhance the truthfulness of an LLM’s responses by modifying hidden state outputs along multiple orthogonal truthful directions identified by a linear probe.
@@
Sparse Autoencoder. The study of sparse autoencoders (SAEs) was first introduced by Agarwal et al. (2016) as an unsupervised learning method for feature extraction. More recently, SAEs have been used to disentangle representations within LLMs for analysing polysemanticity phenomena in these models (Bricken et al., 2023). For example, Cunningham et al. (2023) trained SAEs with L1 loss to map the internal activations of LLMs into a higher-dimensional feature space, using these feature representations to interpret model behaviors. More recently, gated sparse autoencoder (Rajamanoharan et al., 2024) was proposed to address the imbalance between accurate reconstruction and activation sparsity caused by L1 loss biases. Gao et al. (2024) further proposed TopK SAE, leveraging a Top-K activation function to enforce sparsity constraints more effectively to this end.
In this paper, we adopt a TopK SAE to decompose transformer layer representations into a sparsely activated, higher-dimensional feature space, ensuring that the features are decoupled with minimal interference for precise activation steering.
Figure 2: The flowchart of our method. (1) Consistency Relevant Features Locating consists two steps. We first use the method proposed by (Yang et al., 2024) to locate the top-1 transformer layer and then pretrain a SAE and use it to locate the key features responsible for the semantic inconsistencies in the LLM. (2) Steering the LLM towards greater semantic consistency by adjusting the values of the identified key features.
@@ 3 Methodology
As illustrated in Figure 2, our proposed method consists of two main steps: Consistency Relevant Features Locating and Consistency Feature Steering. In the first step, we employ a coarse-to-fine locating strategy. Initially, we use the method proposed by Yang et al. (2024) to identify the top-1 transformer layer that has the most significant impact on the model’s semantic inconsistency. Subsequently, we pretrain a spare autoencoder (SAE) and use it to project the hidden states of the identified transformer layer into a higher-dimensional feature space. Within this feature space, we locate the key features that are most relevant to the LLM’s semantic inconsistency. In the second step, we enable the LLM to produce more semantically consistent responses by adjusting the values of the identified key features.
@@ 3.1 Consistency Relevant Features Locating
Top-1 Transformer Layer Locating. We use the method proposed by Yang et al. (2024) to locate the top-1 transformer layer influences the LLM’s semantic consistency. Specifically, the layer locating dataset $\mathcal{D}_{l}$ was first built, comprising consistent pairs in the form $([m,n],l)$, where $m$ is the input prompt and $n$ is a rephrased version of $m$, which can be generated using large-scale LLMs like GPT-4111https://platform.openai.com/docs/api-reference/chat. The variable $l$ denotes whether the LLM outputs are consistent for $m$ and $n$, which can also be evaluated by large-scale LLMs. Following Yang et al. (2024), the $\mathcal{D}_{l}$ dataset, consisting of 500 instances, was randomly split into a training set and a test set with a 4:1 ratio. Next, for each transformer layer, we train a classifier that takes the concatenated hidden states of the corresponding layer for prompts $m$ and $n$ as inputs and uses the consistency label $l$ as the ground truth.
High classification accuracy on the test set indicates that the layer has a greater influence on the LLM’s semantic inconsistency, while low accuracy suggests that the layer is less relevant to semantic inconsistency. The top-1 performing transformer layer is selected by ranking the test set classification accuracy.
TopK SAE Pretraining. Since the hidden states of the transformer layers contain entangled feature representations (Elhage et al., 2022), this hinders the precise locating of features related to semantic consistency. To address this, it is essential to delve into finer-grained feature-level representations for accurate feature identification.
To achieve this, we pretrain a TopK SAE (Gao et al., 2024) and use it to decompose the dentified top-1 transformer layer representations into a higher-dimensional, sparsely activated feature space, resulting in decoupled feature-level representations.
More specifically, the TopK SAE consists of two main components: the SAE encoder and the SAE decoder. The encoder maps the model’s hidden states into a high-dimensional feature space and applies the TopK activation function to retain the $k$ most significant latent features, ensuring sparse activation. The decoder then reconstructs the original hidden states from these top $k$ latent features. The pretraining process is formally represented as follows:
$\begin{gathered}z(d)=\text{TopK}(W_{\text{enc}}(h_{\text{top-1}}(d)-b_{\text{pre}}))\\ \hat{h}_{\text{top-1}}(d)=W_{\text{dec}}z(d)+b_{\text{pre}},\\ \mathcal{L}=\frac{1}{|\mathcal{D}_{o}|}\sum_{d}\|h_{\text{top-1}}(d)-\hat{h}_{\text{top-1}}(d)\|_{2}^{2},d\in\mathcal{D}_{o}.\\ \end{gathered}$ (1)
@@
Here, $d$ is the data instance from the open domain dataset $\mathcal{D}_{o}$. $h_{\text{top-1}}(d)$ and $\hat{h}_{\text{top-1}}(d)$ denote the hidden states and the reconstructed hidden states from the identified top-1 transformer layer respectively. $W_{\text{enc}}$, $W_{\text{dec}}$, and $b_{\text{pre}}$ represent the encoder weights, decoder weights, and the corresponding bias term, respectively. The $\mathcal{L}$ is the training loss.
@@
Due to the GPU memory limitations, the TopK SAE (Gao et al., 2024) used in our paper decomposes the 4,096 dimensional hidden states to 65,536 dimensional features. We select the top 192 latent features to recover the original hidden states. We leverage 850K public data $\mathcal{D}_{o}$, collected from various domains  [^2] https://huggingface.co/datasets/togethercomputer/RedPajama-Data-1T-Sample to pretrain the TopK SAE. The reason we do not use task-specific data is to ensure that the learned feature representations remain generalizable. The training process is performed with a batch size of 1 per GPU, utilizing 8xV100 GPUs each with 32 GB of memory. We train the TopK SAE for a total of 32,000 steps with the learning rate is set to 2e-4.
@@
Key Features Locating. To identify key features related to semantic inconsistency within the feature space decomposed from the pretrained SAE, we build dataset $\mathcal{D}_{\text{f}}$ consists of 500 contrastive examples in the form $([u,v])$, where $u$ represents the input prompt and $v$ is its rephrased version obtained from GPT-4. The key difference is that $u$ generates a correct result when processed by the LLM, whereas $v$ produces an erroneous response from the target LLM. The correctness labels are assigned based on whether their predicted outputs are identical for NLU tasks. For NLG tasks, GPT-4 is also used to evaluate the correctness label.
@@
Next, we feed the contrastive examples $u$ and $v$ from $D_{\text{f}}$ into the LLM to obtain their respective last token feature representations from the pretrained SAE. We then select key features related to the LLM’s semantic inconsistency by selecting average feature differences from contrastive examples that exceed a specified threshold value $t$. This process is formally represented as follows:
@@
Here, $z(u)$ and $z(v)$ are the last token feature activations of $u$ and $v$ respectively. $g$ is the average feature differences, and $|g|$ is the size of $g$. $\mathcal{I}$ contains the indexes of the located key features. $t$ is a predefined feature difference threshold.
@@ 3.2 Consistency Feature Steering
In this step, we make precise feature representations adjustments to enhance semantic consistency. Specifically, when an input prompt ($q$) is provided during inference, we check whether its activated features indexes from the pretrained TopK SAE are present in the identified key feature indexes $\mathcal{I}$ (See equation 2). If so, we add the corresponding consistency feature bias (average feature differences) on it to guide the LLM towards greater semantic consistency. Formally, this steering process is calculated as follows:
@@
Here, $q$ is the given input prompt, $z(q)$ is the activated features from the pretrained TopK SAE. $\mathds{1}$ denotes the indicator function, which represents the intersection of the located key feature indexes $\mathcal{I}$ and the activated feature indexes of $z(q)$. $b_{i}$ is the consistency feature bias (average feature differences). $\alpha$ is a hyperparameter that controls the strength of steering.